If the position of an object is given by x = 2 m/s2) t3 find a) the average velocity and average acceleration between t=1 and t = 2 and b) the instantaneous velocities and the instantaneous acceleration at t = 1 and t =2 c) compare the average and instantaneous quantities and in each case explain why the one is large

Grade:12th pass

1 Answers

antriksh misri
7 years ago
As x=2t^3So to get v averageWe use Vaverage= total s/total tS travelled at t=1 s is 2Srravelled at t=2 s is 16So total s travelled is 14.•. Total time = 2-1 = 1s.•.v av.=14/1= 14b:-average acceleration=change in velocity/total time elapsedV at t=1s=4*(1)^2=4V at t=2 s=4(2)^2=16Total time=1sec.So av. Acc.=16-4/1=12Inst. Velocity at t=1s is DX/DT=4t^2for t=1sV=4m/s illy for t=2sV=16m/sInst.acc Acc=d2x/dt2=8tFor t=1sInst. Acc. =8For t=2sInst.acc.=16